OpenStudy (anonymous):

Write the verbal sentence as an equation. Then solve the equation. 1.) negative sixteen plus the quotient of a number and 2 is 35. 2.) Thirty-nine minus a number is -19

OpenStudy (anonymous):

1.) -16 +(x/2) = 35==> x/2=51 ==> x=102

OpenStudy (anonymous):

2.) 39-x=-19 ==> x=58